A concrete driveway has a perimeter of 280 feet. The length of the driveway is 20 feet more than 4 times its width. Find the length and width of the driveway by writing two equations with two variables, and solving the equations by either the addition method or the substitution method.
I am having a hard time solving this can some one please help

Answer by checkley77(12569)

LENGTH=4W+20
WIDTH=W
2L+2W=280
2(4W+20)+2W=280
8W+40+2W=280
10W=280-40
10W=240
W=240/10
W=24 ANSWER FOR THE WIDTH.
L=4*24+20
L=96+20=116 ANSWER FOR THE LENGTH.
PROOF
2*24+2*116=280
48+232=280
280=280